# Knowledge Retention for Continual Model-Based Reinforcement Learning

### v1

@@ 1 Introduction
Specifically, we propose DRAGO, a novel continual MBRL approach designed to address catastrophic forgetting and incomplete world models in the absence of prior task data. DRAGO consists of two key components: Synthetic Experience Rehearsal and Regaining Memories Through Exploration. Synthetic Experience Rehearsal uses a continually learned generative model to enable the agent to simulate and learn from synthetic experiences that resemble those from prior tasks. This process allows the agent to synthesize representative transitions that resemble prior experience, reinforcing its understanding of previously learned dynamics without requiring access to past data. In the Regaining Memories Through Exploration component, we introduce an intrinsic reward mechanism that encourages the agent to actively explore states where the previous transition model performs well. This exploration bridges the gap between tasks by discovering connections within the environment, leading to a more comprehensive and cohesive world model. By integrating these two strategies, DRAGO enables the agent to incrementally build a complete understanding of the environment’s dynamics across a sequence of tasks while effectively mitigating catastrophic forgetting. Our empirical results clearly demonstrate that DRAGO achieves superior performance on challenging continual learning scenarios without retaining any data from prior tasks.
@@ 2 Background
In reinforcement learning, an agent interacts with an environment modeled as a Markov Decision Process (MDP). An MDP is defined by a tuple $(\mathcal{S},\mathcal{A},T,r,\gamma)$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $T(s^{\prime}\mid s,a)$ represents the transition dynamics, $r(s,a)$ is the reward function, and $\gamma\in[0,1)$ denotes the discount factor.
@@
3 Dynamics-leaRning while RegAinG MemOries
@@ 3.1 Synthetic Experience Rehearsal
where $\lambda$ is a weighting factor controlling the importance of the synthetic data loss. While this enables the agent to learn from synthetic old experience, the generative model itself (minimizing $-\sum\log p_{G}(\hat{s},\hat{a};\theta)$ in Eqn 4) also requires accumulate the knowledge of different tasks as the training goes on. Retaining such a generative model for every task will also introduces huge additional cost.
@@ 3.3 Overall Algorithm
During training, the dynamics model and the generative model are trained together with the reward&value prediction of the learner and reviewer. At the beginning of each new task, For each new test task, we randomly initialize the reward and value models and reuse only the world model (dynamics).For each new test task, we randomly initialize the reward, policy and value models and reuse only the world model (dynamics).. Moreover, unlike TDMPC, the gradients from updating Q function and reward model are detached for updating the dynamics model in DRAGO. More implementation details can be found in the appendix.
@@ 4.1 Qualitative Results
Figure 4: Prediction score of the learned world models across the entire gridworld after each task. Light color indicates higher prediction accuracy. The heatmaps compare the performance of naive continual training of TDMPC (top row), DRAGO without Synthetic Experience Rehearsal (mid row), with our proposed full DRAGO method (bottom row) after Tasks 1 to 4. The results show that continual MBRL suffers from significant forgetting, maintaining accuracy only in regions relevant to the current task, whereas DRAGO effectively retains knowledge from previous tasks, leading to a more comprehensive world model and improved performance in new tasks.
@@ 4.4 Few-shot Transfer Performance
We also evaluated the agent’s few-shot transfer performance during the continual learning process and compared the results of DRAGO with the other baselines. The setting is useful and common in real-world tasks, especially for robotics, where the number of steps to interact with the environment is limited. Specifically, for each test task in Cheetah and Walker domains, we let the agent train by interacting with the environment for only $20$ episodes and evaluate its average cumulative reward after training. As shown in Table 1, DRAGO outperforms the other baselines in 6 out of 8 tasks. In the two tasks where DRAGO does not outperform, it remains competitive, highlighting its robustness and efficiency in continual learning scenarios.
@@ 5 Related Work
Continual reinforcement learning (CRL) aims to develop agents that can learn from a sequence of tasks, retaining knowledge from previous tasks while efficiently adapting to new ones (Khetarpal et al., 2022; Abel et al., 2023; Anand & Precup, 2023; Baker et al., 2023). Many recent papers investigate the plasticity loss in continual learning (Lyle et al., 2023; Abbas et al., 2023; Dohare et al., 2024). This paper focuses more on how we better retain and aggregate knowledge learned from previous tasks in Continual MBRL, which is related to another central challenge in CRL, catastrophic forgetting, where learning new tasks causes the agent’s performance on earlier tasks to degrade due to the overwriting of important knowledge (McCloskey & Cohen, 1989). To address catastrophic forgetting, several strategies have been proposed: Regularization-Based Methods (Kirkpatrick et al., 2016; Li & Hoiem, 2016; Zenke et al., 2017; Nguyen et al., 2017; Yu et al., 2020a): these approaches introduce constraints during training to prevent significant changes to parameters important for previous tasks. Elastic Weight Consolidation (EWC) (Kirkpatrick et al., 2016) is a prominent example that uses the Fisher Information Matrix to estimate parameter importance and penalize updates accordingly. However regularization-based methods often struggles in practice, especially in reinforcement learning scenarios, due to challenges in accurately estimating parameter importance and scalability issues with large neural networks (Huszár, 2017; Farquhar & Gal, 2018). Replay-Based methods (Riemer et al., 2019; Rolnick et al., 2019; Oh et al., 2022; Henning et al., 2021; Lampinen et al., 2021): these methods typically assume unbounded storage, which is usually not feasible in practice. Our work is therefore focused on the hardest case — alleviating the catastrophic forgetting problem and learn a complete world model without prior data. Generative replay-based methods (Shin et al., 2017; Triki et al., 2017; Rao et al., 2019) share some similar high-level idea with our work. However, none of them has been applied on MBRL or World Models. In terms of Continual MBRL specifically,  Fu et al. (2022) show that the agent can benefit from a joint world model for adapting to new individual tasks. Similarly,  Nagabandi et al. (2019) propose a meta-learning approach where a dynamics model is trained to adapt quickly to new tasks by learning a prior over models. Hypernetwork-based methods (Huang et al., 2021) have been proposed to minimize forgetting while learning task-specific parameters in the multitask setting. Liu et al. (2024) introduces locality-sensitive sparse encoding to learn world models incrementally in a single task online setting.  Kessler et al. (2023) investigate how different experience replay methods will affect the performance of MBRL. Related work for model-based RL in general can be found in Appendix B.
@@ 6 Conclusion
Acknowledgement
This work was conducted using computational resources and services at the Center for Computation and Visualization, Brown University. The authors would like to thank David Abel, Lijing Yu, and Jingwen Zhang for their help and feedbacks on this project.
@@ A.1 Hyperparameters
Appendix B Additional Related Work
B.1 Model-Based Reinforcement Learning

### v2

@@ 1 Introduction
Specifically, we propose DRAGO, a novel continual MBRL approach designed to address catastrophic forgetting and incomplete world models in the absence of prior task data. DRAGO consists of two key components: Synthetic Experience Rehearsal and Regaining Memories Through Exploration. Synthetic Experience Rehearsal uses a continually learned generative model to enable the agent to simulate and learn from synthetic experiences that resemble those from prior tasks. This process allows the agent to synthesize representative transitions that resemble prior experience, reinforcing its understanding of previously learned dynamics without requiring access to past data. In the Regaining Memories Through Exploration component, we introduce an intrinsic reward mechanism that encourages the agent to actively explore states where the previous transition model performs well. This exploration bridges the gap between tasks by discovering connections within the environment, leading to a more comprehensive and cohesive world model. To sum up, we make the following contributions:
  1. 1.
    We introduce DRAGO, a new approach for continual model-based reinforcement learning that addresses catastrophic forgetting while incrementally learning a world model across sequential tasks without retaining any past data.
  2. 2.
    We propose a generative replay mechanism that synthesizes “old” transitions using a learned generative model alongside a frozen copy of the previously trained world model.
  3. 3.
    We design an intrinsic reward signal that nudges the agent toward revisiting states that the old model explained well—effectively “reconnecting” current experiences with previously learned transitions.
  4. 4.
    Through experiments on MiniGrid and DeepMind Control Suite domains, we show that DRAGO substantially improves knowledge retention compared to standard continual MBRL baselines and achieves higher transfer performance, allowing faster adaptation to entirely new (but related) tasks. Code is available at https://github.com/YixiangSun/drago.
@@ 2 Background
In reinforcement learning, an agent interacts with an environment modeled as a Markov Decision Process (MDP). An MDP is defined by a tuple $<\mathcal{S},\mathcal{A},T,r,\gamma>$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $T(s^{\prime}\mid s,a)$ represents the transition dynamics, $r(s,a)$ is the reward function, and $\gamma\in[0,1)$ denotes the discount factor. Throughout this paper, $T$ is the parametric transition model (the learned dynamics predictor) and $p$ denotes a probability or distribution.
@@
3 DRAGO
@@ 3.1 Synthetic Experience Rehearsal
where $\lambda$ is a weighting factor controlling the importance of the synthetic data loss. While this enables the agent to learn from synthetic old experience, the generative model itself (minimizing $-\sum\log p_{G}(\hat{s},\hat{a};\theta)$ in Eqn 4) also requires accumulating the knowledge of different tasks as the training goes on. Retaining such a generative model for every task will also introduces huge additional cost.
@@ 3.3 Overall Algorithm
During training, the dynamics model and the generative model are trained together with the reward&value prediction of the learner and reviewer. At the beginning of each new task, for each new test task, we randomly initialize the reward, policy and value models and reuse only the world model (dynamics).. Moreover, unlike TDMPC, the gradients from updating Q function and reward model are detached for updating the dynamics model in DRAGO. More implementation details can be found in the appendix.
@@ 4.1 Qualitative Results
Figure 4: DRAGO reduces the catastrophic forgetting from previous tasks, illustrated bt the prediction score of the learned world models across the entire gridworld after each task. Light color indicates higher prediction accuracy. The heatmaps compare the performance of naive continual training of TDMPC (top row), DRAGO without Synthetic Experience Rehearsal (mid row), with our proposed full DRAGO method (bottom row) after Tasks 1 to 4. The results show that continual MBRL suffers from significant forgetting, maintaining accuracy only in regions relevant to the current task, whereas DRAGO effectively retains knowledge from previous tasks.
@@ 4.2 Overall Performance
Figure 6: Comparison of DRAGO with (1) Replay-based MBRL, which stores a bounded replay buffer for past tasks (using the same memory size as DRAGO’s generative model), and (2) Pseudo-rehearsal MBRL (Ketz et al., 2019), which generates old data via a pretrained VAE model, without continual training of the generative model as well as our intrinsic reward mechanism. Each plot shows Minigrid transfer performance (average return vs. environment steps) on a new task after sequentially training on earlier tasks. DRAGO consistently achieves higher or faster-improving returns, suggesting stronger knowledge retention and quicker adaptation to the transfer tasks.
@@ 4.4 Few-shot Transfer Performance
We also evaluated the agent’s few-shot transfer performance during the continual learning process and compared the results of DRAGO with the other baselines. The setting is useful and common in real world tasks, especially for robotics, where the number of steps to interact with the environment is limited. Specifically, for each test task in Cheetah and Walker domains, we let the agent train by interacting with the environment for only $20$ episodes and evaluate its average cumulative reward after training. As shown in Table 1, DRAGO outperforms the other baselines in 6 out of 8 tasks. In the two tasks where DRAGO does not outperform, it remains competitive, highlighting its robustness and efficiency in continual learning scenarios.
@@ 5 Related Work
Continual reinforcement learning (CRL) aims to develop agents that can learn from a sequence of tasks, retaining knowledge from previous tasks while efficiently adapting to new ones (Khetarpal et al., 2022; Abel et al., 2023; Anand & Precup, 2023; Baker2023ADA). Many recent papers investigate the plasticity loss in continual learning (Lyle et al., 2023; Abbas et al., 2023; Dohare et al., 2024). This paper focuses more on how we better retain and aggregate knowledge learned from previous tasks in Continual MBRL, which is related to another central challenge in CRL, catastrophic forgetting, where learning new tasks causes the agent’s performance on earlier tasks to degrade due to the overwriting of important knowledge (McCloskey & Cohen, 1989). To address catastrophic forgetting, several strategies have been proposed: Regularization-Based Methods (Kirkpatrick et al., 2016; Li & Hoiem, 2016; Zenke et al., 2017; Nguyen et al., 2017; Yu et al., 2020a): these approaches introduce constraints during training to prevent significant changes to parameters important for previous tasks. Elastic Weight Consolidation (EWC) (Kirkpatrick et al., 2016) is a prominent example that uses the Fisher Information Matrix to estimate parameter importance and penalize updates accordingly. However regularization-based methods often struggles in practice, especially in reinforcement learning scenarios, due to challenges in accurately estimating parameter importance and scalability issues with large neural networks (Huszár, 2017; Farquhar & Gal, 2018). Replay-Based methods (Riemer et al., 2019; Rolnick et al., 2019; Oh et al., 2022; Henning et al., 2021; Lampinen et al., 2021): these methods typically assume unbounded storage, which is usually not feasible in practice. Our work is therefore focused on the hardest case — alleviating the catastrophic forgetting problem and learn a complete world model without prior data. Generative replay-based methods (Shin et al., 2017; Triki et al., 2017; Rao et al., 2019) share some similar high-level idea with our work. However, none of them has been applied on MBRL or World Models. In terms of Continual MBRL specifically,  Fu et al. (2022) show that the agent can benefit from a joint world model for adapting to new individual tasks. Similarly,  Nagabandi et al. (2019) propose a meta-learning approach where a dynamics model is trained to adapt quickly to new tasks by learning a prior over models. Hypernetwork-based methods (Huang et al., 2021) have been proposed to minimize forgetting while learning task-specific parameters in the multitask setting. Liu et al. (2024) introduces locality-sensitive sparse encoding to learn world models incrementally in a single task online setting.  Kessler et al. (2023) investigate how different experience replay methods will affect the performance of MBRL. Related work for model-based RL in general can be found in Appendix B.
@@ 6 Conclusion
7 Limitations
@@
Acknowledgement
This work was supported by the Office of Naval Research (ONR) under grant #N00014-22-12592. Partial funding was also provided by The Robotics and AI Institute. The authors would like to thank David Abel, Lijing Yu, and Jingwen Zhang for their help and feedbacks on this project. This work was conducted using computational resources and services at the Center for Computation and Visualization, Brown University.
@@ A.1 Hyperparameters
Appendix B Related Work (Model-Based Reinforcement Learning)
@@ Appendix E More Ablation Study Results
Figure 9: Prediction score heatmaps for replay-based method (limited storage). The heatmaps illustrates the accuracy of the dynamics model during continual learning by keeping a fixed part of the previous replay buffer (capped to a similar memory size as our generative model). Compared to DRAGO’s performance in Figure 4, this method does not achieve the same qualitative result.